OpenStudy (anonymous):

can someone help me come up with a two-step equation with all real numbers as a solution and show me all the steps?

OpenStudy (rulnick):

2x - 1 = -1 + 2x 2x = 0 + 2x 0 = 0

OpenStudy (rulnick):

Alternative solution: 2x - 1 = -1 + 2x 2x = 0 + 2x x = x